# Braga, Portugal

### v1

@@ Appendix 0.A Requirements in the courses example
Appendix 0.B One-shot system prompt
Alloy is a formal modelling language where that uses sets (named signatures in Alloy) and relations (sets of tuples, named fields in Alloy) to represent information in a domain. Consider, for example, the following Alloy model.
sig Person {
likes : set Car,
owns : Car $\to$ lone Value
sig Adult in Person {}
abstract sig Car {}
sig Sedan, SUV extends Car {}
sig Value {}
This model declares a set named Person, a subset of Person named Adult, a set name Car, two extensions (disjoint subsets) of Car named Sedan and SUV, and a set named Value. It also declares a binary relation named likes, that associates each Person with the set of Cars it likes, and ternary relation named owns that associates each Person with the set of Cars it owns and the respective value. When declaring a field we can use other multiplicity keywords instead of set: set means there is an arbitrary number of target elements associated with a source element, lone means there is at most one related element, some means there is at least one related element, and one means there is exactly one related element. When extensions are declared, if the parent signature is declared as abstract then it can only contain elements that are also contained in its extensions. In this model we have two disjoint extensions of abstract set Car, meaning all cars are either sedans or SUVs. A model can use the module ordering to impose a total ordering on a signature. In this example Value is a totally ordered signature. The total ordering is defined in a binary relation called next, that is implicitly declared.
In Alloy we can express requirements about a model using facts specified in relational logic, which is an extension of first-order logic with relational operators. Given a requirement it is useful to have both positive and negative instances to serve as test cases for validation. An instance of a model is a valuation to all declared sets and relations.
@@ Appendix 0.B One-shot system prompt
For example, if a requirement is "Every person owns a car" the following command specifies a positive instance where this requirement is satisfied.
some disj Person1,Person2,Person3 : Person |
some disj Car1, Car2 : Car |
some disj Value1,Value2,Value3 : Value {
Person = Person1 + Person2 + Person3
Adult = none
Car = Car1 + Car2
Sedan = Car1
SUV = Car2
Value = Value1 + Value2 + Value3
[…]
} for 3 Person, 2 Car, 3 Value expect 1
In the above instance we have three people, named Person1, Person2, and Person3, none of them is an Adult. We have two cars, named Car1 and Car2, the former a Sedan and the latter a SUV. We also have three possible Values, being Value1 the lowest possible and Value3 the highest possible. Person1 owns Car1 with value Value1, and Person2 and Person3 together own Car2 with value Value2. No one likes any car, because relation likes is empty.
In the following interaction you will act as an Alloy expert. I will ask you to generate positive and negative instances for different requirements for a given model.
Here are some instructions:
  • Do not attempt to formalize the requirement with a fact, just output the requested instances as run commands.
  • I will clearly state how many positive and negative instances you should output.
  • All instances should be truly different (be aware that the names of the elements are actually irrelevant).
  • Try to produce minimal instances with few elements.
  • All instances must include a comment explaining in natural language why it is positive or negative.
  • All positive instances should have an expect 1 and all negative an expect 0.
  • All instances must define the scopes for all top-level signatures in the model.
  • In every instance you must specify the values for all declared sets and relations with an equality restriction.
  • To declare the elements of each signature use a some quantifier with the disj keyword to ensure that all elements are different.
  • The value of a relation should take into account its arity. For example, if it is a ternary relation its value must be a set of triples.
  • If a set or relation is empty you must explicitly state that it is equal to the empty set (none) or empty relation (of the correct arity).
  • If there are two relations with the same name you must disambiguate using the <: operator with the domain signature on the left.
  • If a model uses ordering on a signature you must also specify the value of the next binary relation using the <: operator to disambiguate. Signatures with ordering cannot be empty so next is never empty.
  • If the value of some sets and relations is irrelevant for a given requirement you can just state that they are empty or assign them random values.
  • Generate only Alloy code, do not include any explanations outside the code.

### v2

@@ 3.2 Benchmark
For illustration purposes, consider the courses domain model already presented in the previous section. As we have seen, this example uses a few advanced Alloy features, namely non-binary fields, overloading, and the util/ordering module. It contains 15 specification challenges, whose natural language description can be seen in Table 2. The table also shows the number of (semantically distinct) wrong specifications of each requirement, which vary between 33 for requirement “Only professors can teach courses”, and 312 for requirement “A professor cannot teach colleagues”, with an average of 137 per requirement. A higher number of wrong specifications can be a sign of both complexity and ambiguity. However, although some requirements may seem ambiguous, we argue that the one that is codified in the hidden oracle is always the most “correct” interpretation. Take, for example, the requirement with the most number of wrong specifications, “A professor cannot teach colleagues”. At first, many students seem to interpret “colleagues” as any other professor. Under the previous requirements (which should be assumed to be true), that would forbid students that are also professors to be enrolled in any course. Since professors that are also students should be able to enroll, the reasonable interpretation of “colleagues” in this context is the one specified in the oracle: professors that teach some course together. The three other domain models are the following:
  • A photo sharing social network model, where some users are influencers and some photos are ads. This is a simple example in terms of Alloy features (e.g., all fields are binary). However, it is the example with the highest average of wrong specifications per requirement. This could be either because it is the first example we give to students in our courses, so it is natural that they make more mistakes, or because some of the requirements are perhaps not standard in real social networks, which might confuse students.
3.3 Prompt design
Most LLMs nowadays distinguish between system and user prompts: the former clarify the instructions an LLM should follow in the subsequent interactions (for example, defining roles or formatting guidelines for the responses), while the latter contain task requests or questions to be answered according to the system instructions. We used both prompts in our evaluation. The user prompt is simple, asking for test cases for a given requirement and model. Let $R_{i}$ be the natural language formulation of the $i$-th requirement of a domain model, and $N$ be the number of positive and negative test cases to be generated. The user prompt starts as follows, after which the domain model in Alloy is provided.
for the following model. All instances must also satisfy the
requirements $R_{0}$", $\ldots$, and $R_{i-1}$".
Since we want to measure the effect of including examples, we created three versions of the system prompt: zero-shot, one-shot, and few-shot. The three versions were iterated several times in the GPT-5 console until we reached a satisfactory result. These were then fixed and used for the whole evaluation with all LLMs, which was performed via remote calls to the respective APIs. All three system prompts start with a brief presentation of Alloy (a couple of sentences) and end with a definition of the LLM role and an itemized list of precise instructions. The main differences are the following:
Zero-shot
One-shot
Few-shot
The three variants of the system prompt can be found in the GitHub repository.
3.4 Metrics
Tests
The number of test cases to be generated. We ask the LLM to generate $N$ positive and $N$ negative test cases for each requirement. As such, the full test suite should have $2\times N\times 43$ test cases, since we have 43 requirements in total.
Syntax
Consistent
Previous
Valid
The number of test cases satisfying previous requirements that agree with the specified oracle. To measure this we add the oracle specification as a fact to the model, and check if the run command has the expected outcome (satisfiable/unsatisfiable for positive/negative tests). We also report this metric as a percentage of the total number of generated tests (%).
Cost
For RQ5 we evaluated the effectiveness in terms of the ability of the test suite to actually catch wrong specifications. To make fair comparisons, we focus only on the requirements for which the LLM generates a full test suite, with $2\times N$ valid test cases. In the results for this RQ we present the following metrics:
Complete
Wrong
Missed
Mean %
4 Results
4.1 RQ1: influence of prompt design
To assess the influence of prompt design we asked GPT-5 to generate test suites using the three different prompt versions described above with $N=3$. The results are presented in Table 3. As can be seen, the few-shot prompt clearly outperforms the other two versions, being able to generate 247 valid tests out of the 258 requested (a success rate of 96%). Of the 11 test cases that were not valid, 3 failed the syntax check, 3 failed to satisfy the previous requirements, and 5 did not give the expected result according to the oracle. The 11 invalid test cases were distributed among 7 requirements, but all still had at least 3 valid test cases. The success rate drops to 79% for the one-shot prompt and 46% for the zero-shot. Most of this decrease is due to syntactic problems: among the test cases that are executable, 91% and 86% satisfy the previous requirements and the oracle with one-shot and zero-shot prompts, respectively. A possible explanation is the lack of Alloy test case examples in the training data, since most Alloy publications and educational material do not adopt a test-driven methodology. Interestingly, although the input size is much smaller for prompts with fewer examples, the total cost to generate the test suite is higher. The pricing structure for most commercial LLMs imposes a higher price on output tokens when compared to input tokens (sometimes $10\times$ more), and reasoning tokens are charged as output tokens. Since the output is of comparable size, having less information and examples in the system prompt led to the need for more reasoning tokens, which alone justifies the higher cost.
Table 3: Influence of prompt design
Prompt Tests Syntax Consistent Previous Valid % Cost
Few 258 255 255 252 247 96% \$3.56
One 258 226 208 206 205 79% \$3.69
Zero 258 137 120 119 118 46% \$4.20
Table 4: Effect of non-determinism
Run Tests Syntax Consistent Previous Valid % Cost
1st 258 255 255 252 247 96% \$3.56
2nd 258 256 256 256 251 97% \$3.50
3rd 258 252 252 250 246 95% \$3.61
Finding 1: GPT-5 with a few-shot system prompt is highly effective at generating Alloy test suites for structural requirements expressed in natural language. It is also more cost-effective due to less reasoning effort.
4.2 RQ2: effect of non-determinism
LLMs are inherently non-deterministic, and can give different results for identical inputs. Some allow the user to set the so-called temperature, a parameter that affects the randomness of results, by adjusting the probability distribution of the next output token. However, even with temperature set to 0 (the most deterministic setting) determinism is not guaranteed due to hardware factors like GPU concurrency. GPT-5 in particular does not allow the user to set this parameter, so it is even more relevant to study the impact of non-determinism on the results. To assess that effect, we run GPT-5 three times with the same task ($N=3$), and using the few-shot prompt that proved to be more effective. The results are presented in Table 4 (the 1st run is the Few-shot run presented in Table 3). As can be seen, the overall results do not vary significantly, with a success rate always close to 96%. For 33 of the 43 requirements we always got the same number of valid, albeit possibly different, test cases.
Finding 2: The overall effectiveness of GPT-5 when using a few-shot prompt does not seem to be significantly affected by non-determinism.
4.3 RQ3: LLM comparison
To answer this RQ, we compared the effectiveness of GPT-5, version 2025-08-07, with two other state-of-the-art models, Gemini 2.5 Pro from Google DeepMind and Claude Opus 4.1, version 2025-08-05, from Anthropic. We also compared with two less advanced models, namely GPT-5 Mini, version 2025-08-07, a cost-efficient and faster version of GPT-5, also from OpenAI, and the open-source smaller language model Llama 3.1 8B from Meta. Table 5 presents the results, again with $N=3$ and the few-shot prompt. The results in GPT-5 are the ones corresponding to the Few-shot run presented in Table 3. The results from Llama are omitted because this model performed quite badly, and was only able to generate a handful of syntactically correct test cases. Further research is needed to determine if small language models in general are unable to properly cope with this task, or if this result is specific to Llama. The two competing state-of-the-art models performed worse than GPT-5, although still with a success rate above 75%. Gemini 2.5 Pro struggled more with getting test cases that are executable (for example, forgetting to set the scope for a signature), while Claude Opus 4.1 always succeeded in that aspect but struggled to generate test cases that satisfied previous requirements. The main problem for the cost-efficient GPT-5 Mini was getting the syntax right, a pattern similar to the one obtained with GPT-5 for simpler the prompts.
Table 5: LLM comparison
Model Tests Syntax Consistent Previous Valid % Cost
GPT-5 258 255 255 252 247 96% \$3.56
Gemini 258 248 220 212 210 81% \$2.78
Claude 258 258 258 202 197 76% \$5.55
Mini 258 188 183 179 174 67% \$0.53
Table 6: Wrong specification detection
NN Complete Wrong Missed Mean %
1 41 5587 2204 38.10%
2 34 4623 814 17.02%
3 36 4395 525 9.90%
4 35 5129 446 7.45%
5 35 4816 348 6.43%
4.4 RQ4: characterization of invalid test cases
To answer this RQ we manually analyzed all invalid test cases in the 3 runs of GPT-5 reported in Table 4. In total we had 30 invalid test cases, distributed as follows: 11 were syntactically incorrect, 5 failed to satisfy previous requirements, and 14 did not agree with the oracle.
@@ 4.4 RQ4: characterization of invalid test cases
Finding 4: The syntactic errors made by GPT-5 are due to a peculiar feature of Alloy syntax and easily fixable by some post-processing.
Of the 5 test cases that failed previous requirements, one was because the test case was not fully specified and admitted multiple instances, some of them valid and some invalid. No pattern was detected in the remaining 4, all of them indeed failed in one of the previous requirements that were stated in the task prompt. Strangely, in 3 cases where the generated test cases for $R_{i}$ failed to satisfy a previous requirement $R_{j}$ for $j<i$, the LLM had been able to generate valid test cases directly for $R_{j}$.
Of the 14 test cases that did not agree with the oracle, 13 were negative test cases, which seems to suggest that the LLM struggles more with these. The only positive test case that was invalid was for the requirement “Components cannot be their own parts” in the production line example. The invalid test case had a component A that was a part of component B which in turn was part of A. Essentially, the requirement was interpreted literally, but obviously it is a wrong interpretation, as components cannot physically contain themselves direct or indirectly. The 13 invalid negative test cases were distributed among only 3 requirements: 1 for the requirement “Only students work on projects and projects must have someone working on them” in the courses example, because a student that is also a professor was wrongly assumed to not be able to work on a project; 3 for the requirement “Users can see ads posted by everyone, but only see non ads posted by followed users” in the social network example, due to a misinterpretation of the “can” as a “must” (all test cases had a user that did not see an ad, which was considered to be wrong); and 9 for the requirement “A professor cannot teach colleagues” because, likewise many of our students, it was incorrectly assumed that “colleagues” meant any other professor. Note that all 3 runs of GPT-5 made the last two mistakes, which is another indicator that the results are not significantly affected by non-determinism.
4.5 RQ5: effectiveness at detecting incorrect specifications
Finding 6: GPT-5 generates diverse test cases for most requirements.
4.6 Threats to validity
5 Related work
The work by Pan et al. [16] on the generation of model instances using LLMs shares some similarity to ours, since it focuses on structural aspects. Given an XML description of a UML class diagram and a natural language description of a desired instance, LLMs are used to generate the corresponding XMI instance model. A key difference is that we do not expect natural language definitions of the test cases, but only of a high-level requirement. Moreover, their evaluation focuses only on syntactic correctness. To make the technique feasible with open-source, small LLMs, they use an abstract intermediate representation that is converted to XMI in post-processing. Our approach could employ a similar technique to reduce the rate of syntactic errors in the less advanced LLMs.
In the testing community there is some work on using LLMs to generate test cases from natural language requirements (rather than from code), that relate to our work (see, e.g., [24] for a comprehensive survey). For example, Korraprolu et al. [8] present an empirical study to compare the effectiveness of LLMs at generating test cases for Simulink from natural language requirements. Effectiveness was measured as coverage of the (executable) models implemented afterwards (the correctness of the test cases was not assessed). Test cases are simple input-output pairs of numerical or/and discrete values and generated in an tabular form (later translated manually to executable tests). A similar study was done by Xue et al. [22] for the concrete domain of FinTech software [22]. Our tests can also be seen as input-output pairs with a Boolean output but much more complex input (a graph of connected elements from the domain of discourse). Our tests are also directly executable without manual intervention and our evaluation focused on the correctness of the tests.
6 Conclusion and future work
Our experimental evaluation indicates that GPT-5 with a few-shot prompt is highly effective and consistent at generating test suites from natural language requirements, that can be used for the validation of structural requirements of domain models. Generated test cases are almost always syntactically correct, executable, and comply with the stated requirements. The generated test suite for each requirement is also diverse and is very effective at detecting incorrect specifications made by humans. Competing state-of-the-art LLMs are also effective, but less so, when compared with GPT-5. In the future we plan to experiment with alternative more targeted prompts (for example, task specific system prompts adapted for each domain model) and add a post-processing syntax repair procedure, that combined could increase the effectiveness of the technique with open source small language models, making the technique more accessible, namely in educational contexts. We also plan to evaluate the effectiveness of this approach for generating test suites for validating behavioral requirements, namely requirements to be specified with temporal logic. That would make the technique usable in a wider range of formal methods.
Acknowledgments
This work is funded by national funds through FCT – Fundação para a Ciência e a Tecnologia, I.P., under the support UID/50014/2025 (https://doi.org/10.54499/UID/50014/2025).